A firm is considering two mutually exclusive projects, X and Y, with the following cash flows:

0 1 2 3 4
Project X -$1,000 $110 $280 $430 $750
Project Y -$1,000 $1,000 $100 $55 $50

The projects are equally risky, and their WACC is 8%. What is the MIRR of the project that maximizes shareholder value? Round your answer to two decimal places. Do not round your intermediate calculations.

%
Answer #1

X:

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=110/1.08+280/1.08^2+430/1.08^3+750/1.08^4

=$1234.53

NPV=Present value of inflows-Present value of outflows

=$1234.53-$1000

=$234.53(Approx).

Y:

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=1000/1.08+100/1.08^2+55/1.08^3+50/1.08^4

=$1092.07

NPV=Present value of inflows-Present value of outflows

=$1092.07-$1000

=$92.07(Approx).

Hence X is better having higher NPV.

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

Hence future value of inflows=110*(1.08)^3+280*(1.08)^2+430*(1.08)+750

=1679.56032

MIRR=[future value of inflows/Present value of outflows]^(1/time period)-1

=[1679.56032/1000]^(1/4)-1

which is equal to

=13.84%(Approx).